find two consecutive even numbers whose product is 288
Answer by stanbon(75887)

find two consecutive even numbers whose product is 288
---------------------
Even numbers are always a multiple of two:
-----------------------
1st: 2x
2nd: 2x+2
------------
EQUATION:
2x*(2x+2)= 288
4x^2+4x-288 = 0
x^2+x-72 = 0
(x+9)(x-8)=0
x=-9 or x=8
2x=-18 or 2x=16
2x+2=-16 or 2x+2=18
------------------------
Answers:
-18 and -16
OR
16 and 18
==============
Cheers,
Stan H.